User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Ross

Attended: Tiege, Lucas, Patrick, Atifur

Apologies: Arwin, Jonny, Simin, Katie, Mas, Jamie, Lucia


General Issues

  • Yg said they don't get information about the disk space required for an experiment from the new schedule format (.vex experiments?).
    • Experiment only 60GB over and observers didn't notice heaps in advance since vex files not drudged locally, no summary files.
    • Should go to newsmerd:/vlbobs/lba , make directory called experiment (crds96), put the vex file inside there.
  • Yg also said that sometimes the module spreadsheet is not updated (w.r.t. which modules are assigned to experiments).
    • Warren hadn't updated this for crds96, not for two ho only experiments but presumably because not needed by anyone else.
    • Bring up with Warren.

Handover Issues

  • r4848 (ke) - Can't access maserke vnc, but can ping the machine.
  • crds96 (ke) - received “Warning: error tc -301 continuous cal not enabled for dbbc” for each scan. Fixed continuous cal not enabled error by changing cont_cal=off to cont_cal=on in the setup01 function of the procedure file, based off the previous crds95 procedure file.
    • Patrick noticed this the week before. Because experiment drudged manually in the wrong location, it didn't have ke specific procedures.
    • Version was different to normal, suggesting drudged in a different place.
  • aov024 (ho) - Was getting “ERROR sc -13 setcl: formatter to FS time difference 0.5 seconds or greater ”. Ran sy=run setcl offset into the newsmerd fs window (oprin)
  • r1847 (ke) - noticed coordinates had not changed, checked to see the FS had crashed. Restarted FS and the schdule, missing data from 0729 UT till 08:22 UT. It appears log monitor had also stopped responding.
  • r1847 (yg) - Humidity sensor appears to not be working.
    • Could maybe restart monica weather server???
  • crf107 (ho) - rxp is not providing focus position info, instead it is returning error about connection with remote host lost. Live page says receiver position is correct.
  • crf107 (ho) - cryomon had negative pressure, likely a bug.
  • Sound not working on ops8!!
  • Restarted computer and tried a different set of speakers.
  • Sound card not working
  • Use ops4, mobile phone alert and/or byo usb speakers.
  • What are the ftp links for the vex files, if we can't find the summary files?
/home/www/auscope/opswiki/data/pages/operations/agenda2018_06_28.txt · Last modified: 2018/06/28 03:23 by Ross Turner